OpenStudy (warpedkitten):

If f(x) = √x + 12 and g(x) = 2 √x, what is the value of (f-g)(144)?

OpenStudy (anonymous):

\[f(x)=\sqrt{x}+12\] what is \[f(144)=\sqrt{144}+12\]?

OpenStudy (anonymous):

similarly, what is \[g(144)=2\sqrt{144}\]?

OpenStudy (warpedkitten):

No idea how to do that... @satellite73

OpenStudy (anonymous):

do you know what the square root of 144 is? that is what you need

OpenStudy (warpedkitten):

12?

OpenStudy (anonymous):

not 12? but rather 12 !!

OpenStudy (anonymous):

that makes \[f(144)=\sqrt{144}+12=12+12=?\]

OpenStudy (solomonzelman):

satellite73, don't you first subtract g(x) from the f(x) and then plug in 144 for x?

OpenStudy (anonymous):

you can but no need right? \[f-g(144)=f(144)-g(144)\] i would compute each number, then subtract

OpenStudy (anonymous):

we have \[f(144)=24\] all we need is \(g(144)\)

OpenStudy (solomonzelman):

Alright, not typical but actually hadier in a way. Sorry for bugging then:)

OpenStudy (anonymous):

what number is \[g(144)=2\sqrt{144}\]?